OpenStudy (anonymous):

medal for best answer :):):) a toy rocket was launched from the ground. the function f(x)=-16x^2+128x shows the height of the rocket f(x) in feet, from the ground at time x seconds. what it the axis of symmetry of the graph of f(x)and what does it represent?

OpenStudy (anonymous):

a.)x=8takes 8 seconds to reach maximum height, 8 seconds to fall back to the ground. b.)x=4takes 4 seconds to reach maximum height,4 seconds to fall back to the ground. c.)x=8takes 8 seconds to reach maximum height,16 seconds to fall back to the ground. d.)x=4 takes 4 seconds to reach maximum height, 8 seconds to fall back to the ground.

OpenStudy (anonymous):

@surjithayer

OpenStudy (johnweldon1993):

Alright well we see this is an upside down parabola... lets find the vertex (where it changes direction) \[\large -\frac{b}{2a} = -\frac{128}{2(-16)} = \frac{128}{32} = 4\] so the 'x' coordinate of our vertex is 4... so I'm going to say the rocket would take 4 seconds to get up to maximum height...and then it will fall back down taking 4 seconds How do i know it will only take 4 seconds to get down? Well I would find out the x-intercepts of the parabola by factoring -16x^2 + 128x = 0 factor our like terms -16x(x - 8) = 0 so we have 2 equations -16x = 0 and (x - 8) = 0 well that means our 2 x intercepts are 0 and 8 So if it take 4 seconds to go up to the max height...we only have 4 seconds to fall back down Hence the answer being B

OpenStudy (anonymous):

thanks :)